OpenStudy (anonymous):

WILL MEDAL AND FAN!! A right triangle has legs with lengths 9 cm and 40 cm. What is the length of the hypotenuse?

OpenStudy (anonymous):

41cm is the length of the hypotenuse

OpenStudy (anonymous):

how do you find this

OpenStudy (anonymous):

90-41?

OpenStudy (anonymous):

40 is your C right? or is it your B?

OpenStudy (anonymous):

to find the hypotenuse you will have to do the the pythagorean theorem. a2+b2=9^2+40^2=41